# SLS-BRD: A system-level approach to seeking generalised feedback Nash equilibria

### v2

@@ IV-B Price competition in oligopolistic markets
Figure 7: Market game: Relative distance between updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k+1}}^{p})$, top panels, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, bottom panels, for all $p\in\mathcal{P}$.
@@
The evolution of the game given each player’s actions is displayed Figure 5 for the in-game period $t\in(260,400]$ days, after the policy updates have been interrupted. We compare the performance of the policy profile $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}$ with the evolution obtained through the open-loop operation $u(t)=0$. During this period, we simulate a worst-case fluctuation on the baseline demand for the companies’ product by defining
$w(t)=\begin{cases}(\phantom{-}1,\phantom{-}1,\phantom{-}1,\phantom{-}1)&\text{for}\leavevmode\nobreak\ t\in[265,279];\\ (\phantom{-}1,-1,\phantom{-}1,-1)&\text{for}\leavevmode\nobreak\ t\in[293,307];\\ (-1,\phantom{-}1,-1,\phantom{-}1)&\text{for}\leavevmode\nobreak\ t\in[321,335];\\ (\phantom{-}1,\phantom{-}1,-1,-1)&\text{for}\leavevmode\nobreak\ t\in[349,363];\\ (-1,-1,-1,-1)&\text{for}\leavevmode\nobreak\ t\in[377,391],\end{cases}$
@@
and $w(t)=0$ otherwise. The results show that the policy profile obtained by the SLS-BRD routine allows the companies to efficiently respond to changes in their local demands. In general, the players coordinate price changes to alleviate the deviations from the baseline demand caused by the disturbances. Moreover, the player’s actions satisfy the price-cap constraint $\frac{1}{N_{P}}|\sum_{p\in\mathcal{P}}u^{p}(t)|\leq 0.5$ in all disturbance scenarios. The best performance is observed for companies $p\in\{3,4\}$, whereas $p=1$ behaves noticeably worse than all players (especially for $w(t)=(1,-1,1,-1)$ and $w(t)=(-1,1,-1,1)$, when the open-loop operation seems to attain better results). This highlights the fact that fixed-points obtained through the SLS-BRD routine are not necessarily admissible GFNE, and thus might be unfavourable for a subset of players. Finally, we note that these policies are still not able to completely reject the effect of the extreme disturbances: Achieving zero-offset would require incorporating integral action into the player’s policies.
@@
Figure 8: Market game, $t\in(260,400]$ days: State $\bm{x}$ (top panels) and applied control $\bm{Bu}$ (bottom panels) trajectories for an execution of the game $\mathcal{G}_{\infty}^{\text{LQ}}$. The dashed lines refer to the state-trajectories resulting from an open-loop operation of the market with $u(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$
V Concluding remarks
@@
This work presents the SLS-BRD, an algorithm for generalised feedback Nash equilibrium seeking in $N_{P}$-players non-cooperative games. The method is based on the class of best-response dynamics algorithms for decentralised learning and consists of players updating and announcing a parametrisation of their policies until converging to an equilibrium. Because not updating control actions explicitly, this learning dynamics can be performed alongside the execution of the game. Using results from operator theory, we propose that convergence to a Nash equilibrium is geometric if certain conditions are met: Namely, that the best-response maps are contractive operators. Our approach leverages the System Level Synthesis framework to formulate each player’s best-response map as the solution to robust finite-horizon optimal control problems. This framework benefits the SLS-BRD algorithm also by allowing richer information patterns to be enforced directly at the synthesis level. After the main theoretical and practical aspects are discussed, the algorithm is demonstrated on an exemplary problems on the decentralised control of multi-agent systems.

### v3

@@ 4.2 Price competition in oligopolistic markets
Figure 7: Market game: Relative distance between the local updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p})$, left row, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, right row.
The convergence of the SLS-BRD routine to a fixed-point $\bm{K}^{\star}=\bm{\Phi_{u}}^{\star}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{\star}},\ldots,\bm{\Phi_{u}}^{N_{P}^{\star}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 6. Since the exact fixed-point to which the routine will converge is not known for this problem, we let $\bm{\Phi_{u}}^{\star}\approx\bm{\Phi_{u,k_{f}}}$ for $k_{f}=1080$ (when the policy updates are interrupted) and analyse the convergence with respect to this point. In this case, the iterates approach the fixed-point mostly at a geometric rate with the policies requiring roughly $1000$ updates (or $250$ days in-game) before changes become numerically negligible. We stress that the best-responses $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ cannot be (globally) contractive in this case, as the set of fixed points $\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ is not a singleton. Finally, this experiment imply that multi-agent markets might require a half-year of learning dynamics (if policies are updated every $\Delta t=1/4$ [days]) before converging to an equilibrium.
In Figure 7 (left), the relative distances between the individual updates $\{\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p}\}_{p\in\mathcal{P}}$ are displayed. As in the previous example, the updates are of similar magnitude for all players $p\in\mathcal{P}$ and they become numerically negligible at a faster rate than the global convergence in Figure 6. The convergence of the distance between updates is shown to slow considerably, especially for $p\in\{1,4\}$, around $k=50$. Thereafter, these relative distances continue to decrease at a steady rate. In Figure 7 (right), the relative distances $\|\bm{\Phi_{x,k}}^{p}-\bm{\Phi_{x,k}}\|_{\ell_{2}}/\|\bm{\Phi_{x,k}}^{p}\|_{\ell_{2}}$, $k\in\mathbb{N}_{+}$, between responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ and $\bm{\Phi_{x,k}}=\bm{F_{\Phi}}\bm{\Phi_{u,k}}$ are shown. As before, these distances decrease at a similar rate and are relatively small since the initial stages of the game.
@@
The evolution of the game given each player’s actions is displayed Figure 8 for the in-game period $t\in(280,420]$ days, after the policy updates have been interrupted. We compare the performance of the policy profile $\bm{K}^{\star}=\bm{\Phi_{u}}^{\star}\bm{\Phi_{x}}^{-1}$ with the evolution obtained by the open-loop operation $u(t)=0$. During this period, we simulate a worst-case fluctuation on the baseline demand for the companies’ product by defining
$w(t)=\begin{cases}(\phantom{-}1,\phantom{-}1,\phantom{-}1,\phantom{-}1)&\text{for}~{}t\in[285,299];\\ (\phantom{-}1,-1,\phantom{-}1,-1)&\text{for}~{}t\in[313,327];\\ (-1,\phantom{-}1,-1,\phantom{-}1)&\text{for}~{}t\in[341,355];\\ (\phantom{-}1,\phantom{-}1,-1,-1)&\text{for}~{}t\in[369,383];\\ (-1,-1,-1,-1)&\text{for}~{}t\in[397,411],\end{cases}$
@@
and $w(t)=0$ otherwise. The results show that the policy profile obtained by the SLS-BRD routine allows the companies to efficiently respond to changes in their local demands. In general, the players coordinate price changes to alleviate the deviations from the baseline demand caused by the disturbances, while still satisfying the price-cap constraint $\frac{1}{N_{P}}|\sum_{p\in\mathcal{P}}u^{p}(t)|\leq 0.5$. The best performance is observed for the companies $p\in\{3,4\}$, whereas $p=1$ behaves noticeably worse than all players (especially for $w(t)=(1,-1,1,-1)$ and $w(t)=(-1,1,-1,1)$, when the open-loop operation attains better results). This highlights the fact that fixed-points obtained through the SLS-BRD routine are not necessarily admissible GFNE, and might be unfavourable for a subset of players. Finally, we note that these policies are still not able to completely reject the effect of the noise: Achieving zero-offset requires incorporating integral action into the feedback policies.
@@
Figure 8: Market game, $t\in(280,420]$ days: State $\bm{x}$ (top panels) and applied control $\bm{Bu}$ (bottom panels) trajectories for an execution of the game $\mathcal{G}_{\infty}^{\text{LQ}}$. The dashed lines refer to the state-trajectories resulting from an open-loop operation of the market with $u(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$.
5 Concluding remarks
@@
This work presents the SLS-BRD, an algorithm for generalised feedback Nash equilibrium seeking in $N_{P}$-player noncooperative games. The method is based on the best-response dynamics class of algorithms for Nash equilibrium seeking and consists of players updating and announcing a parametrisation of their policies until converging to a fixed-point. Our approach leverages the System Level Synthesis framework to formulate each player’s best-response map as the solution to robust finite-horizon optimal control problems. Because not updating control actions explicitly, this learning dynamics can be performed alongside the execution of the game. This framework also benefits the SLS-BRD by allowing richer information patterns to be enforced directly at the synthesis level. Using results from operator theory, we then established sufficient conditions for this procedure to converge to a generalised feedback Nash equilibrium. After the theoretical aspects are discussed, the algorithm is showcased on exemplary problems on the noncooperative control of multi-agent systems.